OpenStudy (anonymous):

I need help... If f'(x) = 2x/(x^2+1), then f(x) could be... It's a multiple choice answer but i want to know how to get the antiderivative.

OpenStudy (anonymous):

substituion must be used to simplift the integral. let u = x^2 + 1 du = 2x dx this is then equivalent to du/u integral is ln|u| + C which is ln|x^2 + 1| + C

OpenStudy (anonymous):

simplify*

OpenStudy (anonymous):

opposite of differentiation is called as integration \[ f'(x)={2x\over x^2+1}\\ \implies\\ f(x)=\int {2x\over x^2+1}dx\\ \text{solve by substitution:}\quad u=x^2+1\implies du=2xdx \]

OpenStudy (anonymous):

alright i think i'm understanding both of you let me see what f(x) could be using substitution.

OpenStudy (anonymous):

so pretty much like Euler said you get \[\ln |x^{2}+1| +C\] but, there are two answers that are similar but not what i got. \[(A) 3+\ln (x ^{2}+1)\] and \[(D) x+\ln (x^2+1) +C\]

OpenStudy (anonymous):

did they provide an initial condition?

OpenStudy (anonymous):

no initial. pretty much that question or problem i typed up is word for word.

OpenStudy (anonymous):

...other than the multilple choices

OpenStudy (anonymous):

do you want me to list all of the multiple choice answers?

OpenStudy (anonymous):

then the answer is what you have got. C can be any real constant.

OpenStudy (anonymous):

so A then?

OpenStudy (anonymous):

that sounds logical

OpenStudy (anonymous):

alright thank you.